# Continual Learning in Medical Image Analysis: A Comprehensive Review of Recent Advancements and Future Prospects

### v3

@@ 4.4 Hybrid techniques
Combining two or more of individual CL strategies has recently gained interest owing to the enhanced performance gained by harnessing the merits of different CL strategies. Ozdemir et al. (2018) mitigate catastrophic forgetting for incremental medical segmentation applications using a distillation loss inspired by LwF Li and Hoiem (2017). Further, they extend the approach by incorporating memory-replay strategies and task-specific segmentation heads. Specifically, representative images were selected based on abstraction layer response and content distance in the last embedding layer. Liu et al. (2022) also use knowledge distillation and a memory buffer to store the prototypical representation of different organ categories. Akundi and Sivaswamy (2022) also keep class-specific classification head along with knowledge distillation to avoid forgetting for chest x-ray classification application where data arrives in class incremental fashion. A pseudo-rehearsal technique was combined with task-specific dense layers for pneumothorax classification Ravishankar et al. (2019) by Ravishankar et al. (2019); however, it suffers from the linear increase of parameters and unbounded memory to store features for every domain.
Li et al. (2020) introduce dual distillation as well as a fixed memory-based experience replay to continually learn the effective model. Specifically, they maintain an expert classifier, which is nothing but the previous model fine-tuned on new classes after replacing the old fully-connected layer with a new one. Then, a final classifier is distilled using the old classifier and the expert classifier. A fixed amount of old samples is also stored and replayed with current data while training the updated classifier. Then, the nearest-mean-of-exemplars (often calculated by averaging the feature vectors of the stored or selected subset of data in the feature space for each class) method is used for category prediction at the test time.
@@ 4.5 Comparative studies
There have been multiple comparative studies of different kinds of CL approaches on various applications like histopathology-based cancer detection Bándi et al. (2023), histopathology-based tumor detection Kaustaban et al. (2022), chest x-ray classification Lenga et al. (2020), dermatological image modality classification Morgado et al. (2021), etc. Lenga et al. (2020) show a comparative study of joint training and CL techniques such as EWC, and LWF to improve model adaptation to new chest X-ray domains arising from cross-sites while mitigating catastrophic forgetting effects. Upon model evaluation on two datasets from different sites (ChestX-ray14 and MIMIC-CXR) referring to domain shifts, the best results are achieved with joint training, but EWC and LWF offer practical solutions as the previous samples need not be stored with the CL approaches. In the digital pathology research field, Kaustaban et al. (2022) systematically evaluate various CL methods, including regularization-based and rehearsal-based approaches on self-augmented domain shift on the H&E dataset. The authors concluded that though regularization-based methods performed well for DIS and Data-incremental scenarios, only the rehearsal-based method (iCaRL) is effective for CIS which is a more challenging scenario. Further, TIS may be even more challenging for digital pathology as compared to other image domains.
@@
Figure 7: A demonstration for rigid and non-rigid task/domain boundaries for a sequence of four domains in a CL pipeline
@@ 6 Evaluation strategy and metrics
In CL, there is a given sequence of episodes ($1,...,T$) to learn sequentially. Typically, we have well-defined training and test pairs as $(Tr_{i},Te_{i})$ for each $i^{th}$ episode. Upon sequential training, a train-test performance matrix $P\in\mathit{P}^{T\times T}$ is generated for the given sequence of episodes ($1,...,T$). An example matrix for $T=4$ is shown in Figure 8. The cell $P_{t,i}$ is performance on test data of $i^{th}$ episode when the model training up to $t^{th}$ episodes is complete. Thus, after training up to $t^{th}$ episodes, the performance $P_{t,i}$ for $t<i$, $t=i$, and $t>i$ correspond to past, current, and future performance, respectively.
The performance metric is chosen depending on the application. For example, IoU, dice score/coefficient, structural similarity, Hausdorff distance, average symmetric surface distance, etc, are popular choices for segmentation, whereas accuracy, recall, F1-score, and AUC are usually computed for classification applications. Once the performance metric is chosen, we can compute metrics for measuring stability and plasticity on top of it. Various metrics have been proposed to quantify stability (Table 11) and plasticity (Table 12); however, all the metrics are derived from the same train-test performance matrix (Figure 8) to measure the amount of forgetting or forward transfer.
@@ 6.1 Stability: Backward transfer
It refers to measuring the catastrophic forgetting of a model upon learning new episodes. A model is regarded as stable if it gives a similar performance for a particular episode over time. However, upon learning the new episode naively, it is possible to forget the past learning and hence disrupt the performance of the previous episode. In contrast, a carefully designed CL strategy will cause less forgetting or improvement on past episodes upon learning the new episodes. A backward transfer (BWT) metric is used to measure the amount of forgetting. In other words, it is a way to measure stability, i.e., how well the model would retain the previously acquired knowledge to prevent catastrophic forgetting.
Lopez-Paz and Ranzato (2017) define BWT as the influence of new learning $t_{i}$ task on a previous task. When learning a task $t_{i}$ increases the performance on a future task, it is $+ve$ BWT, and if it causes deterioration, it results in $-ve$ BWT. Large $-ve$ BWT is also termed as catastrophic forgetting. The BWT metric given by Lopez-Paz and Ranzato (2017) is very frequently adopted Srivastava et al. (2021); Hofmanninger et al. (2020a) and it is computed after finishing all the episodes as shown in Eq.(A) in Table 11. Average forgetting Derakhshani et al. (2021) is another popularly used metric Derakhshani et al. (2022); Chen and Tang (2022). It is computed as the average of the difference in highest performance, and the final performance reached after training on all episodes is finished (Eq.(B) in Table 11). Chee et al. (2023) compute BWT after learning each subsequent task ($t^{th}$) rather than computing it at the last episode (Eq.(B) in Table 11) as shown through Eq.(C) in Table 11. González et al. (2022) compute BWT as the change in performance after training with each subsequent task averaged over the number of tasks (Eq.(D) in Table 11). Note that BWT is not defined for the last task. Another BWT metric contribution is by Özgün et al. (2020) as shown via Eq.(E) in Table 11.
@@ 6.3 Episode performance and others
Table 13: Literature for CL based classification
@@
Reference (year) Imaging modality Classification of CL scenario CL strategy Model Learning type Evaluation metric Dataset
@@
Ravishankar et al. (2019) Ultrasound, X-ray Cardiac view, Pneumothorax DIS, TIS, IIS Rehearsal + architectural CNN Supervised Accuracy, BWT Subset of ChestXRay (Wang et al. 2017)
@@
Hofmanninger et al. (2020a) CT Chest DIS Rehearsal Res-Net50 (He et al. 2016) Supervised Accuracy, BWT, FWT Private
@@
Lenga et al. (2020) X-ray Chest DIS Comparitive: EWC, LWF DenseNet121 (Huang et al. 2017) Supervised AUC, BWT, FWT ChestX-ray14, MIMIC-CXR
@@
Li et al. (2020) Dermoscopic Skin disease CIS Dual-distillation+rehearsal ResNet18, AlexNet, VGG19 Supervised Accuracy Skin8, Skin40 from ISIC2019 (Tschandl et al. 2018), CIFAR100
@@
Srivastava et al. (2021) X-ray Chest DIS Rehearsal ResNet101 Supervised AUC, FWT, BWT NIH Chest-X-rays14 (McDermott et al. 2020), PadChest (Bustos et al. 2020), CheXpert (Wang et al. 2017)
@@
Bayasi et al. (2021) Dermoscopic Skin lesion DIS Architectural (fixed but partitioned network) ResNet-50 Supervised Accuracy HAM10000 (Tschandl et al. 2018), Dermofit (Ballerini et al. 2013), Derm7pt (Kawahara et al. 2018), MSK (Codella et al. 2018), PH2 (Mendonça et al. 2013), UDA (Codella et al. 2018)
@@
Morgado et al. (2021) Dermoscopic Dermatological imaging modality DIS Regularization, rehearsal VGG-16, MobileNetV2 Supervised Accuracy, precision, recall, F1-score, FWT, BWT Private
@@
Shu et al. (2022) Fundus images Eyes DIS Rehearsal + regularization Resnet-18, 2-layer MLP (Saha et al. 2021) Supervised F1-score, BWT, FWT ODIR, RIADD, REFUGE, iSee (Fang et al. 2020), EyeQ (Fu et al. 2019), PMNIST (Rupesh Kumar et al. 2013)
@@
Akundi and Sivaswamy (2022) X-ray Chest CIS Regularization DenseNet121 Supervised AUC, BWT CheXpert
@@
Kaustaban et al. (2022) Histopathology Tumor (breast cancer, colorectal cancer) IIS, DIS, CIS, TIS Regularization, rehearsal (comparative :EWC, LwF, AGEM, CoPE, iCaRL) ResNet-18 Supervised Accuracy, FWT, BWT CRC (Kather et al. 2019), PatchCam (Bejnordi et al. 2017)
@@
Chen and Tang (2022) Histopathology Breast cancer CIS Regularization Alexnet, Densenet201, Resnet152 Supervised Accuracy, forgetting BreakHis
@@
Derakhshani et al. (2022) Multiple Multiple disease TIS, CIS, cross-DIS Regularization, rehearsal (comparative: EWC, MAS, LwF, iCaRL, EEIL) ResNet18 Supervised Accuracy, forgetting MedMNIST (Yang et al. 2023a)
@@
Chee et al. (2023) Histopathology, retinal, skin lesions Colorectal cancer, eye, skin CIS Rehearsal + architecture + regularization ResNet18 Supervised Accuracy, forgetting CCH5000 (Kather et al. 2016), EyePACS (kag ), HAM10000 (Tschandl et al. 2018)
@@
Zhang et al. (2023c) Multiple Disease (multiple organ source) CIS Rehearsal + architectural ResNet18 Supervised MCR Skin8, Path16, CIFAR100
@@
Bándi et al. (2023) Histopathology Cancer (multi-organ: breast, colon, head-neck) DIS Comparative: (Architecture: PackNet) (Regularization: EWC, GEM) DenseNet Supervised FROC, ROC, Cohen’s Kappa CAMELYON16, CAMELYON17, private for colon and head-neck
@@
Sadafi et al. (2023) Microscopic WBC DIS, CIS, DIS + CIS Rehearsal + regularization ResNeXt-50 Supervised Accuracy Matek-19 (Matek et al. 2019), INT-20, and Acevedo-20 (Acevedo et al. 2020)
Yang et al. (2023b) Multiple Disease classification DIS+CIS Rehearsal ResNet18 Few-shot Accuracy, performance dropping rate MedMNIST (Yang et al. 2023a) (PathMNIST, DermaMNIST, OrganAMNIST, RetinaMNIST, BreastMNIST, BloodMNIST)
@@
Reference (year) Imaging modality Classification of CL scenario CL strategy Model Learning type Evaluation metric Dataset
@@
Baweja et al. (2018) MRI Normal brain structures, white matter lesions TIS Regularization: EWC DeepMedic 3D convnet (Kamnitsas et al. 2017) Supervised DSC coefficient UK Biobank (Miller et al. 2016)
@@
McClure et al. (2018) SMRI Axial and sagittal brain DIS Regularization: DWC MeshNet (Fedorov et al. 2017) (CNN-based) Supervised DSC HCP (Van Essen et al. 2013), NKI (Nooner et al. 2012), Buckner (Biswal et al. 2010), WU120 (Power et al. 2017)
@@
Ozdemir et al. (2018) MRI Humerus and scapula CIS Regularization (LwF) + rehearsal U-Net Supervised DSC, SSD Private
@@
Ozdemir and Goksel (2019) MRI Knee CIS Regularization U-Net Supervised DSC, MSD SKI10 MICCAI Grand Challenge (Heimann et al. 2010)
@@
van Garderen et al. (2019) 4 MR sequences: pre- and post-contrast T1w, T2w, & T2w FLAIR Brain tumor DIS Regularization: EWC 3D U-Net CNN Supervised Dice score BraTS 2018 (Bakas et al. 2017; Menze et al. 2014), private
@@
Özgün et al. (2020) MRI Brain DIS Regularization: MAS QuickNAT (Roy et al. 2019) Supervised CL-DSC, REM, BWT+, TL, FWT CANDI (Kennedy et al. 2012), MALC (Asman and Landman 2013), and ADNI (Jack Jr et al. 2008)
@@
Memmel et al. (2021) MRI Hippocampal DIS U-Net Supervised IoU, DSC (Simpson et al. ), (Kulaga-Yoskovitz et al. 2015), (Boccardi et al. 2015)
@@
Perkonigg et al. (2021) MRI Brain DIS Rehearsal 3D-ModelGenesis (Zhou et al. 2021b) Supervised MAE, FWT, BWT IXI (IXI ), OASIS3 (LaMontagne et al. 2019)
@@
Ranem et al. (2022) MRI Hippocampus DIS Comparative (rehearsal, regularization, hybrid) ViT U-Net Supervised DSC, FWT, BWT Decathlon (Antonelli et al. 2022), Drayd (Denovellis et al. 2021), HarP (Boccardi et al. 2015)
@@
Liu et al. (2022) CT Liver, spleen, pancreas, right kidney left kidney CIS Rehearsal + regularization nnUNet Supervised DC, HD95 (Mul ; Simpson et al. 2019), KiTS (Heller et al. 2019) + private
@@
Karthik et al. (2022) MRI Multiple sclerosis lesions (brain) DIS Rehearsal 3D U-Net Supervised DSC, BWT (Kerbrat et al. 2020)
@@
González et al. (2022) MRI Hippocampus DIS Architectural nnUNet (Isensee et al. 2021) Supervised DSC, FWT, BWT HarP (Boccardi et al. 2015), Dryad (Kulaga-Yoskovitz et al. 2015), Decathlon (Simpson et al. 2019)
Li et al. (2022) CMR Cardiac DIS Rehearsal U-Net Supervised DSC, HD95, BWT, FWT M&Ms (Campello et al. 2021)
@@
Zhang et al.; Zhang et al. (2023d; 2021) MRI Prostate, optic cup and disc DIS Regularization U-Net Supervised DSC, ASSD RUNMC (Bloch et al. 2015), BMC (Bloch et al. 2015), HCRUDB (Lemaître et al. 2015), UCL (Litjens et al. 2014), BIDMC (Litjens et al. 2014), HK (Litjens et al. 2014), (Sivaswamy et al. 2015; Fumero et al. 2011; Orlando et al. 2020)

### v4

@@ 4.4 Hybrid techniques
Combining two or more individual CL strategies has recently gained interest owing to the enhanced performance gained by harnessing the merits of different CL strategies. Ozdemir et al. (2018) mitigate catastrophic forgetting for incremental medical segmentation applications using a distillation loss inspired by LwF Li and Hoiem (2017). Further, they extend the approach by incorporating memory-replay strategies and task-specific segmentation heads. Specifically, representative images were selected based on abstraction layer response and content distance in the last embedding layer. Liu et al. (2022) also use knowledge distillation and a memory buffer to store the prototypical representation of different organ categories. Akundi and Sivaswamy (2022) also keep class-specific classification head along with knowledge distillation to avoid forgetting for chest X-ray classification application where data arrives in class incremental fashion. A pseudo-rehearsal technique was combined with task-specific dense layers for pneumothorax classification Ravishankar et al. (2019) by Ravishankar et al. (2019); however, it suffers from the linear increase of parameters and unbounded memory to store features for every domain.
Li et al. (2020a) introduce dual distillation as well as a fixed memory-based experience replay to continually learn the effective model. Specifically, they maintain an expert classifier, which is nothing but the previous model fine-tuned on new classes after replacing the old fully-connected layer with a new one. Then, a final classifier is distilled using the old classifier and the expert classifier. A fixed amount of old samples is also stored and replayed with current data while training the updated classifier. Then, the nearest-mean-of-exemplars (often calculated by averaging the feature vectors of the stored or selected subset of data in the feature space for each class) method is used for category prediction at the test time.
@@ 4.5 Comparative studies
There have been multiple comparative studies of different kinds of CL approaches on various applications like histopathology-based cancer detection Bándi et al. (2023), histopathology-based tumor detection Kaustaban et al. (2022), chest X-ray classification Lenga et al. (2020), dermatological image modality classification Morgado et al. (2021), etc. Lenga et al. (2020) show a comparative study of joint training and CL techniques such as EWC, and LwF to improve model adaptation to new chest X-ray domains arising from cross-sites while mitigating catastrophic forgetting effects. Upon model evaluation on two datasets from different sites (ChestX-ray14 and MIMIC-CXR) referring to domain shifts, the best results are achieved with joint training, but EWC and LwF offer practical solutions as the previous samples need not be stored with the CL approaches. In the digital pathology research field, Kaustaban et al. (2022) systematically evaluate various CL methods, including regularization-based and rehearsal-based approaches on self-augmented domain shift on the H&E dataset. The authors concluded that though regularization-based methods performed well for DIS and IIS, only the rehearsal-based method (iCaRL) is effective for CIS which is a more challenging scenario. Further, TIS may be even more challenging for digital pathology as compared to other image domains.
@@
Figure 7: A demonstration for rigid and non-rigid task/domain boundaries for an example sequence of four domains in a CL pipeline
@@ 6 Evaluation strategy and metrics
In CL, there is a given sequence of episodes ($1,...,T$) to learn sequentially. Typically, we have well-defined training and test pairs as $(Tr_{i},Te_{i})$ for each $i^{th}$ episode. Upon sequential training, a train-test performance matrix $P\in\mathit{P}^{T\times T}$ is generated for the given sequence of episodes ($1,...,T$). An example matrix for $T=4$ is shown in Figure 8. The cell $P_{t,i}$ is performance on test data of $i^{th}$ episode when the model training up to $t^{th}$ episodes is complete. Thus, after training up to $t^{th}$ episodes, the performance $P_{t,i}$ for $i<t$, $i=t$, and $i>t$ correspond to performances on past, current, and future episodes, respectively.
The performance metric is chosen depending on the application. For example, IoU, dice score/coefficient, structural similarity, Hausdorff distance, average symmetric surface distance, etc., are popular choices for segmentation, whereas accuracy, recall, F1-score, and AUC are usually computed for classification applications. Once the performance metric is chosen, we can compute metrics for measuring stability and plasticity on top of it. Various metrics have been proposed to quantify stability (Table 15) and plasticity (Table 16); however, all the metrics are derived from the same train-test performance matrix (Figure 8) to measure the amount of forgetting or forward transfer.
@@ 6.1 Stability: Backward transfer
It refers to measuring the catastrophic forgetting of a model upon learning new episodes. A model is regarded as stable if it gives a similar performance for an already learnt episode over time. However, upon learning the new episode naively, it is possible to forget the past learning and hence disrupt the performance of the previous episode. In contrast, a carefully designed CL strategy will cause less forgetting or improvement on past episodes upon learning the new episodes. A backward transfer (BWT) metric is used to measure the amount of forgetting. In other words, it is a way to measure stability, i.e., how well the model would retain the previously acquired knowledge to prevent catastrophic forgetting.
Lopez-Paz and Ranzato (2017) define BWT as the influence of new learning $t_{i}$ task on a previous task. When learning a task $t_{i}$ increases the performance on a previous task, it is $+ve$ BWT, and if it causes deterioration, it results in $-ve$ BWT. Large $-ve$ BWT is also termed as catastrophic forgetting. The BWT metric given by Lopez-Paz and Ranzato (2017) is very frequently adopted Srivastava et al. (2021); Hofmanninger et al. (2020a) and it is computed after finishing all the episodes as shown in Eq.(A) in Table 15. Average forgetting Derakhshani et al. (2021) is another popularly used metric Derakhshani et al. (2022); Chen and Tang (2022). It is computed as the average of the difference in highest performance, and the final performance reached after training on all episodes is finished (Eq.(B) in Table 15). Chee et al. (2023) compute BWT after learning each subsequent task ($t^{th}$) rather than computing it at the last episode (Eq.(B) in Table 15) as shown through Eq.(C) in Table 15. González et al. (2022) compute BWT as the change in performance after training with each subsequent task averaged over the number of tasks (Eq.(D) in Table 15). Note that BWT is not defined for the last task. Another BWT metric contribution is by Özgün et al. (2020) as shown via Eq.(E) in Table 15.
@@ 6.3 Episode performance and others
Table 17: Literature for CL based classification. (**) indicates non-imaging modality
@@
Reference (year) Imaging modality Application CL scenario CL strategy Model Learning type Evaluation metric Dataset
@@
Ravishankar et al. (2019) X-ray, Ultrasound (a) Cardiac view classification, (b) Pneumothorax identification DIS, TIS, IIS Rehearsal + Architectural CNN Supervised Accuracy, BWT Subset of ChestXRay (Wang et al. 2017)
@@
Hofmanninger et al. (2020a) CT Classification of synthetic object in chest images DIS Rehearsal Res-Net50 (He et al. 2016) Supervised Accuracy, BWT, FWT Private
@@
Lenga et al. (2020) X-ray Chest X-ray classification DIS Comparitive: EWC, LWF DenseNet121 (Huang et al. 2017) Supervised AUC, BWT, FWT ChestX-ray14, MIMIC-CXR
@@
Li et al. (2020a) Dermoscopic Skin disease classification CIS Regularization + Rehearsal ResNet18, AlexNet, VGG19 Supervised Accuracy Skin8, Skin40 from ISIC2019 (Tschandl et al. 2018), CIFAR100
@@
Srivastava et al. (2021) X-ray Chest X-ray classification DIS Rehearsal ResNet101 Supervised AUC, FWT, BWT NIH Chest-X-rays14 (McDermott et al. 2020), PadChest (Bustos et al. 2020), CheXpert (Wang et al. 2017)
@@
Bayasi et al. (2021) Dermoscopic Skin lesion classification DIS Architectural (fixed but partitioned network) ResNet-50 Supervised Accuracy HAM10000 (Tschandl et al. 2018), Dermofit (Ballerini et al. 2013), Derm7pt (Kawahara et al. 2018), MSK (Codella et al. 2018), PH2 (Mendonça et al. 2013), UDA (Codella et al. 2018)
@@
Morgado et al. (2021) Dermoscopic Dermatological imaging modality classification DIS Regularization, rehearsal VGG-16, MobileNetV2 Supervised Accuracy, precision, recall, F1-score, FWT, BWT Private
@@
Shu et al. (2022) Fundus images Eyes disease classification DIS Rehearsal + Regularization Resnet-18, 2-layer MLP (Saha et al. 2021) Supervised F1-score, BWT, FWT ODIR, RIADD, REFUGE, iSee (Fang et al. 2020), EyeQ (Fu et al. 2019), PMNIST (Rupesh Kumar et al. 2013)
@@
Akundi and Sivaswamy (2022) X-ray Chest X-ray classification CIS Regularization DenseNet121 Supervised AUC, BWT CheXpert
@@
Kaustaban et al. (2022) Histopathology Tumor classification (breast cancer, colorectal cancer) IIS, DIS, CIS, TIS Regularization, Rehearsal (comparative:EWC, LwF, AGEM, CoPE, iCaRL) ResNet-18 Supervised Accuracy, FWT, BWT CRC (Kather et al. 2019), PatchCam (Bejnordi et al. 2017)
@@
Chen and Tang (2022) Histopathology Breast cancer identification CIS Regularization Alexnet, Densenet201, Resnet152 Supervised Accuracy, forgetting BreakHis
@@
Derakhshani et al. (2022) Multiple Multiple disease classification TIS, CIS, cross-DIS Regularization, Rehearsal (comparative: EWC, MAS, LwF, iCaRL, EEIL) ResNet18 Supervised Accuracy, forgetting MedMNIST (Yang et al. 2023a)
@@
Chee et al. (2023) Histopathology, retinal, skin lesions (a) Colorectal cancer detection, (b) diabetic retinopathy classification, (c) skin lesion classification CIS Rehearsal + Architecture + Regularization ResNet18 Supervised Accuracy, forgetting CCH5000 (Kather et al. 2016), EyePACS (kag ), HAM10000 (Tschandl et al. 2018)
@@
Zhang et al. (2023c) Multiple Disease classification (multiple organ source) CIS Rehearsal + Architectural ResNet18 Supervised MCR Skin8, Path16, CIFAR100
@@
Bándi et al. (2023) Histopathology Cancer detection cross organs (breast, colon, head-neck) DIS Comparative: (Architecture: PackNet) (Regularization: EWC, GEM) DenseNet Supervised FROC, ROC, Cohen’s Kappa CAMELYON16, CAMELYON17, private for colon and head-neck
@@
Sadafi et al. (2023) Microscopic WBC classification DIS, CIS, DIS + CIS Rehearsal + Regularization ResNeXt-50 Supervised Accuracy Matek-19 (Matek et al. 2019), INT-20, and Acevedo-20 (Acevedo et al. 2020)
Yang et al. (2023c) Multiple Disease classification DIS+CIS Rehearsal ResNet18 Few-shot Accuracy, performance dropping rate MedMNIST (Yang et al. 2023a) (PathMNIST, DermaMNIST, OrganAMNIST, RetinaMNIST, BreastMNIST, BloodMNIST)
Table 18: Literature for CL based classification. (**) indicates non-imaging modality (Continued)
Reference (year) Imaging modality Application CL    scenario CL strategy Model Learning type Evaluation metric Dataset
Byun et al. (2023) 2D fundus, dermoscopic (a) Diabetic retinopathy severity classification, (b) skin lesion classification DIS Rehearsal ViT-B/16 supervised AUC, backward transfer fundus (Messidor-2, APTOS), skin lesion (BCN2000, PAD-UEFS-20, HAM10000)
Xiao et al. (2023) dermoscopic & clinical images Skin disease classification CIS Rehearsal (Memory buffer optimization) ResNet-18 Few-shot accuracy, sensitivity, and specificity private curated from various public sources (ISIC-2018, ISIC-2019, ISIC-2020, SD-198, SD-256, PAD-OFED-20, Asan, Hally, etc.)
Yang et al. (2023b) Dermoscopic, clinical Skin disease classification CIS Architectural ResNet101 + GMMs Supervised MCR Skin7 (Codella et al. 2018), Skin40 (Sun et al. 2016)
Verma et al. (2023) (a) Fundus, (b) pathology Disease classification CIS Comparative (Rehearsal, Regularization, Architecture) ResNet50 Supervised MCR (a) OCT (Kermany et al. 2018), (b) PathMNIST (Yang et al. 2023a)
Huang et al. (2023) WSI Tumor subtype classification CIS Rehearsal ConvNeXt (image encoder) + transformer (classifier) supervised BWT, AUC, Masked AUC, Accuracy 4 datasets (NSCLC, BRCA, RCC, ESCA) from TCGA [^2] https://www.cancer.gov/ccg/research/genome-sequencing/tcga project
Hua et al. (2023) sEMG** Gesture classification CIS Rehearsal+ Regularization CNN supervised Accuracy, Remember Ninapro DB2 (Atzori et al. 2014)
Sun et al. (2023b) time-series signals** Mortality identification, Sepsis identification CIS Rehearsal LSTM online BWT, FWT, AUC, Accuracy COVID-19 (Yan et al. 2020) , SEPSIS (Seymour et al. 2017)
Sun et al. (2023a) time-series signals** Multi-class classification CIS Rehearsal + Regularization 1D-ConvNet few-shot Accuracy, relative performance drop rate Mit-BIH (Goldberger et al. 2000), FaceAll (Dau et al. 2019), UWave (Dau et al. 2019), Mit-BIH Long-Term ECG (Goldberger et al. 2000)
Li and Jha (2023) Physiological signals** Disease classification DIS, CIS, TIS Rehearsal MLP Supervised Accuracy, F1 score, BWT CovidDeep(Hassantabar et al. 2021), DiabDeep(Yin et al. 2019), MHDeep(Hassantabar et al. 2022)
Kim et al. (2024) ECG** Arrhythmia detection DIS Comparative (Regularization: LwF, EWC, MAS) 1D-CNN+MLP Supervised AUC (Zheng et al. 2020), (Wagner et al. 2020), (Alday et al. 2020), (Liu et al. 2018)
Aslam et al. (2024) Time series signal** Disease classification DIS Regularization (EWC) LSTM Supervised Accuracy, forgetting Mpox (Mathieu et al. 2022), Influenza (cdc ), and Measles (eur )
Ceccon et al. (2024b) X-ray Chest disease classification NCI Rehearsal + Regularization Supervised AUC, F1 score, forgetting ChestX-ray14(Wang et al. 2017), CheXpert(Irvin et al. 2019)
Verma et al. (2024) (a) Fundus, (b) histopathology Disease classification TIS Architectural ResNet50 Supervised Accuracy (a) OCT (Kermany et al. 2018), (b) PathMNIST (Yang et al. 2023a)
Bayasi et al. (2024) (a) Dermoscopic, (b) microscopic, (c) histopathology (a) Skin lesion classification, (b) blood cell classification, (c) Colon tissue classification CIS, DIS, CIS+DIS Regularization + Architectural ResNet50 Supervised Avg. recall, forgetting, AUPRC (a) HAM10000 (Tschandl et al. 2018), Dermofit (Ballerini et al. 2013), Derm7pt (Kawahara et al. 2018), MSK (Codella et al. 2018), UDA (Codella et al. 2018), BCN (Combalia et al. 2019), PH2 (Mendonça et al. 2013) (b) PBS-HCB (Acevedo et al. 2020), (c) NCT-CRC-HE (Kather et al. 2019)
Ceccon et al. (2024a) X-ray Chest disease classification CIS Comparative (Regularization, Rehearsal, hybrid) ResNet50 Supervised AUC, TPR gap ChestX-ray14 (Wang et al. 2017), CheXpert (Irvin et al. 2019)
Thandiackal et al. (2024) Histopathology Tissue classification DIS Rehearsal ResNet18 Unsupervised F1 score K-19 (Kather et al. 2016), K-16 (Kather et al. 2019), CRC-TP (Javed et al. 2020)
Qazi et al. (2024) (a) Histopathology, (b) dermoscopic Disease classification CIS Architectural ResNet18 Supervised Accuracy, FLOPS Skin8, Path16
Bringas et al. (2024) Time-series (motion-sensor)** Alzheimer’s disease stage identification IIS (2 Ep.,3 Ep.,4 Ep.) Rehearsal 1D-CNN Supervised Accuracy, F1-score, forgetting private
@@
Reference (year) Imaging modality Application CL scenario CL strategy Model Learning type Evaluation metric Dataset
@@
Baweja et al. (2018) MRI Brain structure segmentation (normal, white matter lesions) TIS Regularization: EWC DeepMedic 3D convnet (Kamnitsas et al. 2017) Supervised DSC coefficient UK Biobank (Miller et al. 2016)
@@
McClure et al. (2018) SMRI Brain structure segmentation (axial and sagittal) DIS Regularization: DWC MeshNet (Fedorov et al. 2017) (CNN-based) Supervised DSC HCP (Van Essen et al. 2013), NKI (Nooner et al. 2012), Buckner (Biswal et al. 2010), WU120 (Power et al. 2017)
@@
Ozdemir et al. (2018) MRI Segmentation of humerus and scapula CIS Regularization (LwF) + rehearsal U-Net Supervised DSC, SSD Private
@@
Ozdemir and Goksel (2019) MRI Knee segmentation CIS Regularization U-Net Supervised DSC, MSD SKI10 MICCAI Grand Challenge (Heimann et al. 2010)
@@
van Garderen et al. (2019) 4 MR sequences: pre- and post-contrast T1w, T2w, & T2w FLAIR Brain tumor segmentation DIS Regularization: EWC 3D U-Net CNN Supervised Dice score BraTS 2018 (Bakas et al. 2017; Menze et al. 2014), private
@@
Özgün et al. (2020) MRI Brain MRI segmentation DIS Regularization: MAS QuickNAT (Roy et al. 2019) Supervised CL-DSC, REM, BWT+, TL, FWT CANDI (Kennedy et al. 2012), MALC (Asman and Landman 2013), and ADNI (Jack Jr et al. 2008)
@@
Memmel et al. (2021) MRI Hippocampal segmentation DIS U-Net Supervised IoU, DSC (Simpson et al. ), (Kulaga-Yoskovitz et al. 2015), (Boccardi et al. 2015)
@@
Perkonigg et al. (2021) MRI Brain MRI segmentation DIS Rehearsal 3D-ModelGenesis (Zhou et al. 2021b) Supervised MAE, FWT, BWT IXI (IXI ), OASIS3 (LaMontagne et al. 2019)
@@
Ranem et al. (2022) MRI Hippocampus segmentation DIS Comparative (rehearsal, regularization, hybrid) ViT U-Net Supervised DSC, FWT, BWT Decathlon (Antonelli et al. 2022), Drayd (Denovellis et al. 2021), HarP (Boccardi et al. 2015)
@@
Liu et al. (2022) CT Organ segmentation (liver, spleen, pancreas, right kidney left kidney) CIS Rehearsal + regularization nnUNet Supervised DC, HD95 (Mul ; Simpson et al. 2019), KiTS (Heller et al. 2019) + private
@@
Karthik et al. (2022) MRI Multiple sclerosis lesions (brain) segmentation DIS Rehearsal 3D U-Net Supervised DSC, BWT (Kerbrat et al. 2020)
@@
González et al. (2022) MRI Hippocampus segmentation DIS Architectural nnUNet (Isensee et al. 2021) Supervised DSC, FWT, BWT HarP (Boccardi et al. 2015), Dryad (Kulaga-Yoskovitz et al. 2015), Decathlon (Simpson et al. 2019)
Li et al. (2022) CMR Cardiac segmentation DIS Rehearsal U-Net Supervised DSC, HD95, BWT, FWT M&Ms (Campello et al. 2021)
@@
Zhang et al.; Zhang et al. (2023d; 2021) MRI (a) Prostate segmentation, (b) optic cup and disc segmentation DIS Regularization U-Net Supervised DSC, ASSD RUNMC (Bloch et al. 2015), BMC (Bloch et al. 2015), HCRUDB (Lemaître et al. 2015), UCL (Litjens et al. 2014), BIDMC (Litjens et al. 2014), HK (Litjens et al. 2014), (Sivaswamy et al. 2015; Fumero et al. 2011; Orlando et al. 2020)